eelyaK11:

The work that Ryan did to find the greatest common factor of 48 and 72 is shown below. Prime factorization of 48: 2, 2, 2, 2, 3 Prime factorization of 72: 2, 2, 2, 3, 3 The greatest common factor is 2 ´ 2 ´ 2 ´ 2 ´ 3 ´ 3. What is Ryan’s error? Ryan did not include all the common factors in the greatest common factor. Ryan did not list all the factors in the prime factorization of 48. Ryan did not list all the factors in the prime factorization of 72. Ryan found the least common multiple.

JUNIOR42:

@Five

JUNIOR42:

@MrMudd183

eelyaK11:

@Five

eelyaK11:

@Five

Five:

@supie

eelyaK11:

huh u dont know

supie:

Obviously not.

eelyaK11:

lol help me pls

supie:

Ok, sure! I'll brb tho.

eelyaK11:

ok

Five:

i was helping my brother

supie:

Ok, so the easy way to figure it out would be to do \[72-48=?\] So the, count by ? which is _ ,_, _

supie:

Does that make sense?

supie:

@eelyaK11??

supie:

Oh ok.

eelyaK11:

oof

eelyaK11:

no i didnt my parent called my name

Five:

y you give me medal supie

Five:

oh wow

eelyaK11:

help needed

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Ok, so the easy way to figure it out would be to do \[72-48=?\] So the, count by ? which is _ ,_, _ \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Does that make sense? \(\color{#0cbb34}{\text{End of Quote}}\)

eelyaK11:

24

Five:

24 ok whats nxt

supie:

\[24 + 24\] or \[24(2)\]

eelyaK11:

ok

eelyaK11:

48

supie:

now \[48+24=?\]

eelyaK11:

72

Five:

correct

supie:

Ok, so the easy way to figure it out would be to do \[72-48=24\] So the, count by ? which is 24 ,48, 72 So now, what do you think the answer is @eelyak11?

eelyaK11:

uhh gimmie a sec

eelyaK11:

a?

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @eelyaK11 a? \(\color{#0cbb34}{\text{End of Quote}}\) No. Why A?

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Ok, so the easy way to figure it out would be to do \[72-48=24\] So the, count by ? which is 24 ,48, 72 So now, what do you think the answer is @eelyak11? \(\color{#0cbb34}{\text{End of Quote}}\) This, what we just solved was the LCM I believe.

eelyaK11:

oohhhhh

eelyaK11:

then it would hve to be D

supie:

Correct!

eelyaK11:

yes can you stick around help me with few more

supie:

Sure!

eelyaK11:

i made a tiny list

supie:

k, but make a new post.